# Towards Large-Scale In-Context Reinforcement Learning by Meta-Training in Randomized Worlds

### v2

@@ Abstract
In-Context Reinforcement Learning (ICRL) enables agents to learn automatically and on-the-fly from their interactive experiences. However, a major challenge in scaling up ICRL is the lack of scalable task collections. To address this, we propose the procedurally generated tabular Markov Decision Processes, named AnyMDP. Through a carefully designed randomization process, AnyMDP is capable of generating high-quality tasks on a large scale while maintaining relatively low structural biases. To facilitate efficient meta-training at scale, we further introduce step-wise supervision and induce prior information in the ICRL framework. Our results demonstrate that, with a sufficiently large scale of AnyMDP tasks, the proposed model can generalize to tasks that were not considered in the training set. The scalable task set provided by AnyMDP also enables a more thorough empirical investigation of the relationship between data distribution and ICRL performance. We further show that the generalization of ICRL potentially comes at the cost of increased task diversity and longer adaptation periods. This finding carries critical implications for scaling robust ICRL capabilities, highlighting the necessity of diverse and extensive task design, and prioritizing asymptotic performance over few-shot adaptation.
**footnotetext: Equal Contributionfootnotetext: Corresponding to: wangfan@cuhk.edu.cn, shaoshanliu@cuhk.edu.cn
@@ 1 Introduction
Large-scale pre-training has achieved tremendous success, especially in processing natural languages, images, and videos [1, 2, 3, 4]. They have demonstrated the ability to address unseen tasks through in-context learning (ICL) [5], a paradigm that leverages contextual information to enhance performance. Unlike in-weight learning (IWL), which relies on gradient-based updates to model weights, ICL enables models to acquire new skills in a few-shot manner, thereby enhancing their adaptability to novel environments. Sharing commonalities with model-based meta-learning approaches [6, 7], ICL can accommodate traditional learning paradigms within its framework, including supervised learning [7, 8], imitation learning [9, 10, 11], and reinforcement learning [12, 13, 14, 15]. This significantly alleviates the need for laborious human-designed objective functions and optimization strategies, which are typically required in IWL. Further, gradient-based IWL has been criticized for its inefficiency in continuously adapting to new tasks [16]. In contrast, ICL has demonstrated plasticity that resembles the adaptability of the human brain [17].
On the one hand, pre-training on massive uncurated data faces limitations in terms of the controllability of ICL. The mechanisms underlying the emergence of ICL capabilities and their limitations are not yet fully understood [18]. Additionally, the robustness of ICL in pre-training is under question [19]. On the other hand, meta-learning aims to develop in-context learning functionality rather than master specific skills. However, the efficacy of these approaches is frequently constrained by the scarcity of large, diverse datasets and the narrow scope of available tasks—limitations particularly pronounced in meta-reinforcement learning (MetaRL) and in-context reinforcement learning (ICRL). Consequently, most existing studies evaluate a model’s adaptive capacity solely within narrowly defined domains, thereby restricting the broader generalization of these methods.
To further enhance the scalability and generality of ICRL, we extend the widely used multi-armed bandit benchmark [20] to tabular Markov Decision Processes (MDPs). We introduce AnyMDP, a scalable task generation framework where MDPs are designed with fully randomized transition and reward functions. To produce non-trivial environments with minimal structural bias, we propose a procedural generation method that synthesizes diverse tasks, creating significant challenges for learners. To scale ICRL efficiently, we further enhance the modeling and training framework with two key innovations: step-wise supervision and prior knowledge induction. This allows us to meta-train models on an unprecedented scale—over 6 billion time steps and context lengths exceeding 512K. The model, referred to as OmniRL, was trained exclusively on AnyMDP tasks and demonstrated robust generalization to entirely unseen Gymnasium environments. This serves to validate the efficacy of the framework. Using the scalable task set, we further investigate how data distribution impacts ICRL performance at scale. Our experiments reveal three critical insights: (1) data coverage and extensiveness are crucial for effective meta-learning, as these factors delineate the boundaries between IWL-dominated task identification and ICL-dominated general-purpose learning paradigms; and (2) asymptotic performance (i.e., long-term adaptation) can be a more reliable indicator of ICRL capability than zero-shot or few-shot performance metrics; (3) there is considerable potential for training regimes that prioritize task diversity and extensiveness, over fidelity to prior experiences.
To summarize, our contributions includes: (1) We introduce AnyMDP, a scalable tabular MDP generation framework designed to support the training of ICRL on over $100K$ tasks across tens of billions of simulation steps. In addition, we propose an efficient training and model framework for ICRL that ensures scalability. (2) We examine how data distribution affects ICRL and empirically verify the need for diverse and extensive task design, along with the importance of evaluating asymptotic performance.
@@ 2.1 Emergence of In-Context Learning
Meta-learning, also known as learning to learn [21], pertains to a category of approaches that prioritize the acquisition of generalizable adaptation skills across a spectrum of tasks. It encompasses a broad array of methodologies, including gradient-based optimization [22] and model-based meta-learning [6]. Meta-learning concepts align with the formation of System 3 [23], which involves synaptic recalibration for environmental adaptation and complementing LLMs’ System 1 (intuitive thinking) and System 2 (in-context reasoning). Large models pre-trained with huge, uncurated datasets also incentivize ICL in a manner similar to meta-learning [5, 9, 24, 25, 26, 27, 28]. For clarity, "pre-training" describes training targeting skill acquisition, typically followed by gradient-based tuning, while "meta-training" refers to training for acquiring learning ability without requiring subsequent gradient-based tuning. Numerous theoretical investigations [19, 29, 30, 31] have been conducted to rigorously uncover the association between ICL ability and pre-training data distribution. Empirical evidence demonstrates that burstiness significantly enhances ICL performance in both language modeling tasks [30] and decision-making problems [26]. Current ICL implementations predominantly function as few-shot learners [32, 33], particularly in scenarios with insufficient task diversity. Analyses and experiments have been conducted to show that computation-based ICL can exhibit a richer array of behaviors than gradient-based IWL [34, 35, 36], particularly in terms of plasticity and continual learning [17], which makes ICL potentially a generalized learner. However, recent studies indicate that models tend to prioritize reliance on IWL over ICL whenever feasible [31]. Our research aligns with the growing body of work on general-purpose in-context learning [32, 37, 38], which emphasizes ICL as the primary mechanism for skill acquisition rather than IWL.
@@ 2.2 In-Context Reinforcement Learning
In-context reinforcement learning (ICRL) encompasses algorithms that dynamically adapt to contextual information by synthesizing self-generated trajectories and incorporating external feedback [6, 12]. It typically employs recurrent neural structures [6] and attention structures [12, 20] that are capable of encoding the historical interactions in the inner loop. The meta-training process that optimizes the parameters specifying the learning mechanisms of the inner loop is called outer loop. Common choices for the outer-loop optimizer for ICRL include reinforcement learning [6, 13, 20], evolutionary strategies [39, 40], and supervised learning [12, 15]. While supervised learning generally achieves greater sample efficiency compared to RL and evolutionary strategies, it often suffers from weaker regret bounds [41, 42]. Additionally, the absence of an oracle policy—a key requirement for supervised learning—can pose a critical bottleneck. In such cases, frameworks like RL Coaches [9, 43] offer viable alternatives by bridging the gap between supervised paradigms and reinforcement learning. Another significant challenge in the study of In-context Reinforcement Learning (ICRL) stems from the scarcity of large-scale task sets, as task diversity and scale are critical for ICL ability [44]. Existing research predominantly employs domain randomization techniques to expand canonical reinforcement learning environments into broader classes of related environments. Building general-purpose adaptive agents in an open-ended world has attracted much attention in recent years [37, 45, 46, 47, 48]. However, existing environments suffer from significant structural biases, which hinder the direct applicability of ICRL models trained on one task class to another.
2.3 Domain Randomization
Domain Randomization (DR) was originally introduced to improve sim-to-real transfer [49]. In Meta-learning, it is also common to randomize a subset of domain parameters to create these variant tasks. These benchmarks can be typically categorized into randomizing the rewards or targets while keeping the transitions fixed [12, 15, 20, 22, 50, 51], randomizing the dynamics while keeping the targets unchanged [39, 40, 52, 53, 54], and randomizing the observations and labels without altering the underlying transitions and rewards [32, 55, 56]. In addition to static rule-based randomization, automatic domain randomization [46, 57] introduces a dynamic randomization process by automatically searching for the best performance in the target domain. Although DR creates a class of tasks with a certain variety, it is restricted by the original task setting. Recently, researchers have proposed generating entirely new tasks through randomization, a method designed to minimize structural biases [38, 58, 59]. This approach represents a promising direction for advancing scalable ICRL with minimum inductive biases.
3 Methodology
3.1 Generating AnyMDP tasks by World Randomization
Motivations.To promote general-purpose ICL, we argue that a critical step is reducing the shared structure (common ground) across tasks in the training set. By reducing dependence on IWL, the trained model will instead rely primarily on ICL to solve specific tasks. This suggests prioritizing task diversity and coverage over fidelity to real-world scenarios, ensuring the model generalizes beyond narrow patterns. Inspired by these advancements, we introduce world randomization, which extends beyond traditional domain randomization to further diversify the task set. It is designed to minimize inductive biases in task formulation, thereby further enhancing task diversity and coverage.
@@ 3.1 Generating AnyMDP tasks by World Randomization
Formulations. We denote a task of Markov Decision Processes (MDPs) with $\tau=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E},T_{E},\gamma\rangle$ where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ the state-transition probability distribution, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ the reward function, $\mathcal{S}_{0}$ the set of initial states, and $\mathcal{S}_{E}$ the set of terminal states, $T_{E}$ the maximum steps in an episode, and $\gamma$ the discount factor. We consider only fully visible discrete MDPs with $\mathcal{S}=\{1,...n_{s}\}$, $\mathcal{A}=\{1,...n_{a}\}$, denoted by $\tau(n_{s},n_{a})$. With these notations, domain randomization with base task $\tau$ can be represented as $\mathcal{DR}(\tau)=\langle\mathcal{S},\mathcal{A},\mathcal{P}+\epsilon_{P},\mathcal{R}+\epsilon_{R},\mathcal{S}_{0},\mathcal{S}_{E},T_{E}\rangle$.
We then formally define the AnyMDP task set $\mathcal{T}(n_{s},n_{a})$, which is constructed through MDPs and world randomization, with $|\mathcal{S}|\leq n_{s}$ and $|\mathcal{A}|\leq n_{a}$, as:
$\mathcal{T}(n_{s},n_{a})=\{\tau(n_{s},n_{a})|\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E},T_{E}\sim\text{Random}\}$ (1)
with $\mathcal{R}(s,a,s^{\prime})\sim\mathcal{N}(\mu_{\mathcal{R}}(s,a,s^{\prime}),\Sigma_{\mathcal{R}}(s,a,s^{\prime}))$, $\mathcal{P}$, $\mu_{\mathcal{R}}$, and $\Sigma_{\mathcal{R}}$ sampled randomly from $\mathbb{R}^{n_{s}\times n_{a}\times n_{s}}$, $\mathcal{S}_{0},\mathcal{S}_{E}\subset\mathcal{S}$, and $T_{E}\in[T_{min},T_{max}]$.
Procedural Generation Algorithm. Although Equation 1 can be used to represent any tasks specified by MDPs, in practice, there are inevitable biases induced by the sampling methods. The most straight-forward methods generate $\mathcal{P},\mu_{\mathcal{R}},\Sigma_{\mathcal{R}}$ using Gaussian and uniform distributions. However, our analysis revealed that tasks generated in this way exhibited simplistic structural patterns. To effectively search “high-quality” MDPs that present meaningful challenges to learners, we designed the procedural generation algorithm based on the following principle:
Under a uniform random policy, the probability of reaching high-valued states is always over 00 but decreases at least exponentially with $n_{s}$.
We use a banded transition kernel and composite reward function to satisfy this constraint, while preserving lower structural bias in the sampling process. Further analyses of the procedural generation are left to Section B.1. In subsequent sections, we further demonstrate that AnyMDP tasks achieve the broadest coverage to date across task sets defined within discrete state and action spaces.
@@
3.2 Modeling and Training Framework

### v3

@@ Abstract
In-Context Reinforcement Learning (ICRL) enables agents to learn automatically and on-the-fly from their interactive experiences. However, a major challenge in scaling up ICRL is the lack of scalable task collections. To address this, we propose the procedurally generated tabular Markov Decision Processes, named AnyMDP. Through a carefully designed randomization process, AnyMDP is capable of generating high-quality tasks on a large scale while maintaining relatively low structural biases. To facilitate efficient meta-training at scale, we further introduce decoupled policy distillation and induce prior information in the ICRL framework. Our results demonstrate that, with a sufficiently large scale of AnyMDP tasks, the proposed model can generalize to tasks that were not considered in the training set through versatile in-context learning paradigms. The scalable task set provided by AnyMDP also enables a more thorough empirical investigation of the relationship between data distribution and ICRL performance. We further show that the generalization of ICRL potentially comes at the cost of increased task diversity and longer adaptation periods. This finding carries critical implications for scaling robust ICRL capabilities, highlighting the necessity of diverse and extensive task design, and prioritizing asymptotic performance over few-shot adaptation.
[…]
@@ 1 Introduction
In-Context Learning (ICL) [1] has emerged as a pivotal paradigm for large pre-trained models [2, 3, 4, 5], enabling adaptation to novel tasks without parameter adjustments. Unlike gradient-based in-weight learning (IWL), which modifies parameter weights through optimization, ICL facilitates skill acquisition through natural interaction. This is achieved by unifying versatile learning paradigms by sequence of interactions in the context, including supervised learning [6, 7], imitation learning [8, 9, 10], and reinforcement learning [11, 12, 13], etc. By leveraging those contexts, ICL eliminates reliance on manually engineered objective functions and the labor-intensive optimization infrastructure required by traditional IWL frameworks, providing versatile, self-directed manners of learning. ICL further aligns with black-box meta-learning principles [14, 11] where task-agnostic adaptation occurs through computation and memorization rather than explicit parameter tuning. It can be regarded as a unifying framework that bridges the gap between conventional meta-learning approaches and contemporary large-scale language model capabilities.
The historical conception of ICL has been predominantly associated with supervised few-shot learning paradigms. Recently, the frameworks for learning from dynamic experiences have achieved growing success, where the contextual information can encompass observations, actions, feedback, and reasoning. Among these emerging paradigms, In-Context Reinforcement Learning (ICRL) stands out as a critical approach. Unlike traditional ICL, which arises from self-supervised pre-training, ICRL typically requires explicit training via reinforcement learning (RL) algorithms [12], introducing substantially higher computational complexity and training costs for large-scale applications. While recent supervised learning approaches have shown promise in enhancing ICRL efficiency [13, 15, 16], current implementations remain constrained by limited context lengths (typically <1K tokens) and narrow task scales (often <1K distinct tasks). These limitations significantly restrict ICRL’s capacity for broad task generalization, highlighting an urgent need for scalable architectures that can reconcile contextual flexibility with computational efficiency.
Additionally, existing benchmarks for ICRL predominantly fall into two categories: 1) oversimplified environments like multi-armed bandit tasks, or 2) procedurally generated seed tasks such as maze navigation [17, 18, 19] or arcade-style games [20]. While the former lacks practical relevance due to their constrained problem spaces, the latter suffers from superficial parameter randomization - diversifying only peripheral elements like visual textures or sub-objectives while preserving core environmental biases (e.g., fixed transition dynamics and reward functions). This design paradigm confines ICRL agents to narrow problem distributions, yielding systems that demonstrate limited generalization scope. Consequently, scaling task complexity beyond a few hundred variants encounters diminishing returns, as models inherit fundamental limitations from their training environments’ structural biases.
To advance the scalability and generalization capabilities of ICRL, this work presents two interrelated contributions: a novel benchmarking task set and a scalable ICRL framework. First, we introduce AnyMDP, a scalable task generation environment where Markov Decision Processes (MDPs) are systematically designed with fully randomized transition dynamics and reward functions. To balance environmental richness with learner challenge, we propose the procedural generation process remarked with banded transition matrices. This design minimizes structural biases while preserving problem complexity, creating a benchmark that demands genuine contextual adaptation rather than exploiting environmental regularities. Second, we propose Decoupled Policy Distillation (DPD) to enhance ICRL training efficiency. By disentangling reference policy from behavior policy, and inducing prior knowledge into the context, we achieve unprecedented training scales and ICRL versatility. Our implementation, OmniRL, was trained exclusively on AnyMDP tasks using context sequences of up to 512K steps per sequence and 6 billion steps overall. Empirical validation demonstrates OmniRL’s capability of many-shot generalization to entirely novel tasks, confirming our models generalization scope and versatile learning ability.
Moreover, leveraging the scalable AnyMDP task set, we conduct ablations that reveal critical insights into ICRL’s dependence on task diversity and long-context modeling. First, the boundary between IWL and ICL is fundamentally determined by training-task coverage: sufficiently diverse distributions (at least 10K unique tasks in our case) elicit emergent task learning capabilities, whereas limited coverage reduces systems to task recognition regimes. Second, long-context modeling proves indispensable for interpreting complex task specifications and achieving broad generalization. These findings establish a new paradigm for ICRL research in which scalable benchmarks and context-aware training jointly yield systems that approach human-like adaptability in dynamic decision-making domains.
@@ 2.1 Emergence of In-Context Learning
Meta-learning, also known as learning to learn [21, 22, 12], pertains to a category of approaches that prioritize the acquisition of generalizable adaptation skills across a spectrum of tasks. Large models pre-trained on massive, uncurated datasets naturally give rise to In-Context Learning (ICL), a phenomenon analogous to model-based meta-learning [1, 8, 23, 24, 25, 26, 27]. Investigations reveal that ICL ability is tightly linked to pre-training data distribution, with factors spanning burstiness and intra-sequence correlation [28, 29, 25] to overall diversity [30]. Complementary analyses show that computation-based ICL can emulate a rich spectrum of learning behaviors, including gradient descent [31, 32, 33], Bayesian inference [33], and broader learning paradigms [13]. Those findings suggest its potential as a general-purpose learning machine [34, 35, 36] rather than a mere mechanism for instruction following or few-shot adaptation. Building on the recent progress in ICL and the principles of meta-learning, we adopt the term meta-training to denote the training procedure that explicitly equips the model with the capacity to learn in context, thereby distinguishing it from pre-training, which yields only zero-shot skills and incentivizes ICL as an incidental by-product of exposure to uncurated data.
@@ 2.2 In-Context Reinforcement Learning
In-context reinforcement learning (ICRL) encompasses algorithms that dynamically adapt to decision-making tasks by synthesizing self-generated trajectories and incorporating external feedback in the context [12, 13]. Unlike one-off ICL adaptation, ICRL stresses continual policy improvement driven by accumulated experience, echoing the “third system” forged through ongoing interaction and experience accumulation, as articulated by Barabasi et al. [37]. ICRL typically employs recurrent  [12, 38] and attention-based [13, 39] neural structures that are capable of encoding the interactive histories in the inner loop. The meta-training process that searches the parameters for learning functionality is called outer loop. Common choices for the outer-loop optimizer for ICRL include reinforcement learning [12, 40, 39], evolutionary strategies [41, 17], and supervised learning [13, 16]. While supervised learning generally achieves higher sample efficiency compared to RL and evolutionary strategies, it often suffers from the bottleneck of distribution shift [42, 43].Additionally, the absence of an oracle policy can become a critical bottleneck for supervised learning; consequently, alternatives replace the expert with an RL coach for data synthesis [8, 44]. A further obstacle to ICRL research is the lack of large-scale task suites, which hampers the emergence of ICRL in real-world decision-making problems and motivates the creation of procedurally generated tasks at scale.
2.3 Procedurally Generated Tasks
Commonly employed procedural-generation techniques include randomizing a subset of domain parameters to create variant tasks by randomizing the rewards or targets while keeping the transitions fixed [13, 16, 39, 22, 45, 46], randomizing the dynamics while keeping the targets unchanged [41, 17, 20, 47, 48, 18], and randomizing the observations and labels without altering the underlying transitions and rewards [34, 19, 49]. These techniques is frequently applied to enhance the robustness of decision and policy model for sim-to-real transfer [50, 51, 52], namely domain randomization (DR). Although DR creates a class of tasks with a certain variety, it is restricted by the original task setting and basically forms a close and finite task set. Recently, generating open-ended tasks with enhanced diversity has emerged as a focal point in research, with a significant emphasis on game-based frameworks [53, 51, 36, 54, 53, 51, 55]. Such approaches are increasingly recognized as vital inductive biases for fostering adaptability. While procedurally generated discrete Markov Decision Processes (MDPs) provide a promising avenue for minimizing structural bias, owing to their configuration via a constrained set of hyperparameters, naive sampling of these MDPs often results in a concentration of trivial tasks, which lack meaningful complexity [56, 57]. In this work, we address this limitation by strategically incorporating critical features during the procedural generation of discrete MDPs.
3 AnyMDP: Procedural Generation of High-Quality MDP Tasks
3.1 Problem Setting and Definitions
@@
We denote a task of Markov Decision Processes (MDPs) with $\tau=\langle\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\mathcal{S}_{0},\mathcal{S}_{E},T_{E},\gamma\rangle$ where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ the state-transition probability distribution, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ the reward function, $\mathcal{S}_{0}$ the set of initial states, and $\mathcal{S}_{E}$ the set of terminal states, $T_{E}$ the maximum steps in an episode, and $\gamma$ the discount factor. We consider only fully visible discrete MDPs with $\mathcal{S}=\{1,...n_{s}\}$, $\mathcal{A}=\{1,...n_{a}\}$, denoted by $\tau$. We define average state transition $\hat{\mathcal{P}}_{\mathfrak{r}}(s^{\prime}|s)=\mathbb{E}_{a\sim\pi^{\mathfrak{r}}(a|s)|}$, which is the transition under uniform policy $\mathfrak{r}:a\sim\pi^{\mathfrak{r}}(a|s)=\frac{1}{|\mathcal{A}|}$. For absorbing states $s\in\mathcal{S}_{E}$, the transition was connected back to $\mathcal{S}_{0}$. We then formally define the AnyMDP task collection $\mathcal{T}(n_{s},n_{a},\eta,\epsilon,\kappa,b_{-},b_{+})$ as MDPs satisfying the following conditions:
  • Ergodic: average state transition $\hat{\mathcal{P}}_{\mathfrak{r}}$ has unique positive stationary distribution.
  • Banded Transition Matrix: there exists a ranking of states: $s_{1},...,s_{n_{s}}$, such that: $\hat{\mathcal{P}}_{\mathfrak{r}}$ satisfies the following condition:
    $\hat{\mathcal{P}}_{\mathfrak{r}}(s_{j}|s_{i})$ $\equiv 0\text{, }\forall j<i-b_{-}\text{, or }j>i-b_{+}$
    $\sum_{j<i}\hat{\mathcal{P}}_{\mathfrak{r}}(s_{j}|s_{i})$ $>\eta\text{, if }i>b_{-},\sum_{j>i}\hat{\mathcal{P}}_{\mathfrak{r}}(s_{j}|s_{i})>\epsilon\text{, if }i<n_{s}-1$ (1)
  • Ascending Value Function: given the aforementioned ranking of states, the value function $V^{*}$ under the optimal policy satisfies the following condition:
    $V^{*}(s_{n_{s}})>\max_{s\in\mathcal{S}_{0}}V^{*}(s)+\kappa$ (2)
@@
3.2 Motivations of AnyMDP
AnyMDP is motivated by the following visions: (1) Existing procedurally generated MDP benchmarks like PROCON [56] and Garnet [57] overlook terminal states ($\mathcal{S}_{E}$) which is a critical feature of real-world environments where agents enter terminal states (e.g., task completion or failure). Therefore, AnyMDP allows terminal states to match many real tasks. (2) Through banded transition matrix and ascending value function, AnyMDP has high-valued states that are exponentially less possible to reach by random exploration. In real-world tasks, especially long-horizon ones, the probability of reaching the goal by naïve random choices decays exponentially with task length. Consequently, solely relying on random exploration can be extremely inefficient in solving those tasks; instead, a continuous exploration–exploitation trade-off is required to approach the optimum gradually. Inspired by this necessity, we impose on the Markov Decision Process the constraint that states with higher values are reached with lower probability, thereby guiding the learner toward systematic, intelligent exploration rather than mere random search.
By defining the stationary distribution $p_{\mathfrak{r}}$ with $p_{\mathfrak{r}}P_{\mathfrak{r}}=p_{\mathfrak{r}}$, the upper and lower bound of the chances of arriving at any states is ensured by the following theorem:
Theorem 1.
Given the condition of Equation˜1, for $j>b$, where $b=\max(b_{-},b_{0})+b_{+}$, there exists a value $0<\delta<1/(b_{+}+1)$. If $\eta>1-\delta$,the sampling algorithm ensures that the transition $\hat{\mathcal{P}}_{\mathfrak{r}}$ has a unique positive stationary distribution (Ergodic). Specifically, for the state $s_{j}$, the $p_{\mathfrak{r}}(s_{j})$ satisfies the following bounds:
$C_{1}\epsilon^{j-b}<p_{\mathfrak{r}}(s_{j})<C_{2}(1-\delta)^{j-b}$,
where $C_{1}$ and $C_{2}$ are positive constants.
Theorem˜1 guarantees that the probability of staying at $s_{j}$ by random choice of actions decreases at least exponentially with respect to $j$, while also being bounded above by another exponentially decaying, yet non-negligible, probability to ensure ergodicity. Using a banded transition matrix, we further equip the MDP with a Composite Reward (CR) that decomposes reward generation into independent components, ensuring an ascending value function while preserving low structural bias when sampling. We have relegated the proof of Theorem˜1 and the details of the sampling procedures to the Appendix˜B for brevity.
3.3 Comparison with Other Procedural MDPs and Empirical Validation
Figure 1: An ablation study comparing AnyMDP tasks with Garnet MDP and AnyMDP without composite reward demonstrates that the procedural generation algorithm of AnyMDP produces tasks of higher learning difficulty.
Figure 2: Comparison of the stationary distributions (SDs) of the oracle policy and the uniform random policy across four classes of environments: AnyMDP - $\tau(16/64,5)$, and Garnet MDP-$(16/64,5,2/4)$ demonstrates that AnyMDP exhibits uniquely exponentially-decaying SDs. The results are averaged over 64 randomly sampled tasks, with the SDs for each task re-ranked in descending order.
The proposed sampler guarantees high-quality MDPs in three aspects: ergodicity, low structural bias, and high learning difficulty. To empirically verify the high learning difficulty and to validate Theorem˜1, we run two groups of experiments. Figure 1 compares the performance of two well-known discrete MDP learning methods: Tabular Q-Learning [59] with Upper Confidence Bound (TQL-UCB) [60] and Proximal Policy Optimization (PPO) [61], on tasks sampled from AnyMDP and Garnet MDP [57]. Garnet uses parameters $n_{s},n_{a},b,\sigma,\tau$ to shape generated MDPs. Here, we set $\sigma=0.1,\tau=0,b=\{2,4\}$, and keeps $n_{s}$ and $n_{a}$ equal to AnyMDP for comparison. To isolate the impact of composite reward (CR) Equation˜10 sampling, we also include a baseline without the composite reward sampling technique (AnyMDP w/o CR). The results show that AnyMDP tasks pose greater challenges for both RL methods. Figure 2 further illustrates the stationary distribution (SD) of states by tasks sampled from AnyMDP and Garnet MDPs. It demonstrates exponential decay in SD of AnyMDP, empirically validating Theorem 1, while Garnet MDPs exhibit flatter SDs, especially under the random policy.
4 The Scalable ICRL Framework of OmniRL